# An operator learning perspective on parameter-to-observable maps

### v1

@@ 5.1. Moments of an advection–diffusion model
Our first model problem concerns a canonical advection–diffusion PDE in two spatial dimensions. This equation often arises in the environmental sciences and is useful for modeling the spread of passive tracers (e.g.: contaminants, pollutants, aerosols), especially when the driving velocity field is coupled to another PDE such as the Navier–Stokes equation. Our setup is as follows. Let $\mathcal{D}=(0,1)^{2}$ be the spatial domain and $\mathsf{n}$ denote the unit inward normal vector to $\mathcal{D}$. For a prescribed time-independent velocity field $v\colon\mathcal{D}\to\mathbb{R}^{2}$, the state $\phi\colon\mathcal{D}\times\mathbb{R}_{>0}\to\mathbb{R}$ solves
@@
for all $x=(x_{1},x_{2})\in\mathcal{D}$. Note that $u$ is constant in the vertical $x_{2}$ direction. The eigenvalues $\{\tau_{j}\}_{j\in\mathbb{N}}$ and eigenfunctions $\{e_{j}\}_{j\in\mathbb{N}}$ correspond to the Mercer decomposition of a Matérn covariance function over $\mathbb{R}$ with smoothness exponent $1.5$ and lengthscale $0.25$ [61]. The eigenfunctions are restricted to $(0,1)\subset\mathbb{R}$. We choose
$z_{j}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathsf{Uniform}([-1,1])\quad\text{for all}\quad j\in[d_{\mathrm{KL}}]\,.$
Thus, up to normalization constants, the velocity field (40) is the (truncated) KL expansion of a subgaussian stochastic process. We take the input to either be the full $x_{1}$-velocity field $u\colon\mathcal{D}\to\mathbb{R}$ or the iid realizations $z\coloneqq(z_{1},\ldots,z_{d_{\mathrm{KL}}})^{\top}$ of the random variables that affinely parametrize $u$.
Velocity Input
State Output
@@
Define the nonlinear quantity-of-interest map $q^{\dagger}\colon L^{4}(\mathcal{D};\mathbb{R})\to\mathbb{R}^{4}$ as follows. First, for any $h\in L^{2}(\mathcal{D};\mathbb{R})$, let
@@
Hence, $q^{\dagger}_{1}$ is the mean, $q^{\dagger}_{2}$ the standard deviation, $q^{\dagger}_{3}$ the skewness, and $q^{\dagger}_{4}$ the excess kurtosis. Our goal is to build FNM surrogates for the PtO map that sends the input representation (either the full velocity field or its finite number of iid coefficients) to the QoI values of the state $\phi$ at final time $t=3/4$ (see Figure 3). Therefore, we train FNMs to approximate each of the following ground truth maps:
@@
The training data is obtained by solving (39) with a second-order Lagrange finite element method on a mesh of size $32\times 32$ and Euler time step $0.01$. For each $d_{\mathrm{KL}}\in\{2,20,1000\}$, we generate $10^{4}$ iid data pairs for training, $1500$ pairs for computing the test error (which is (37) over the $1500$ test pairs instead of over the $N$ training pairs), and $500$ pairs for validation. All FNM models with 2D spatial input or output functions use $12$ modes per dimension and a channel width of $32$. For the V2V-FNM, we use a 1D latent function space with $12$ modes and channel width of $96$. We compare all FNM models to a standard fully-connected NN with $3$ layers and constant hidden width $2048$. These architecture settings were selected based on a hyperparameter search over the validation dataset for $d_{\mathrm{KL}}=1000$ that mimics the parameter complexity experiments in [7, 34]. The models are trained on the relative loss (37) for $500$ epochs in $L^{2}$ output space norm for functions and Euclidean norm for vectors. The optimizer settings include a minibatch size of $20$, weight decay of $10^{-4}$, and an initial learning rate of $10^{-3}$ which is halved every $100$ epochs. We train $5$ iid realizations of the models for various values of $N$ and $d_{\mathrm{KL}}$ and report the results in Figure 4.
@@
Figure 4 reveals several interesting trends. In general, training models to emulate the advection–diffusion PtO map with finite-dimensional vectors as input is more difficult than adopting function space input variants of the problem. The difficulty is further exacerbated as the dimension of the input vector (here, $d_{\mathrm{KL}}$) increases. We hypothesis that this gap in performance would reduce if the vector input models received the weighted KL coefficients $\{\sqrt{\tau_{j}}z_{j}\}$ as input instead of the iid sequence $\{z_{j}\}$. This way the model would have access to decay information and hence an ordering of the coefficients. The standard finite-dimensional NN performs poorly across all KL expansion dimensions. The training of the NN is also quite erratic, as evidenced by the large green shaded regions indicating large variance over multiple training runs. The output space seems to play less of a role than the input space. Indeed, the F2F and F2V FNMs with function space inputs generally achieve the lowest test error regardless of $N$ and $d_{\mathrm{KL}}$. The full-field F2F method slightly out performs the end-to-end F2V method by a small constant factor (except for when $d_{\mathrm{KL}}=2$). Since $q^{\dagger}$ is a smoothing QoI due to its integral definition, this observation aligns with the theoretical insights from Subsection 4.3.2. The fast convergence of some of the FNM models, especially for the low-dimensional cases $d_{\mathrm{KL}}=2$ and $d_{\mathrm{KL}}=20$, could potentially be explained by the lack of noise in the data, the smoothness of the QoIs, and the nonconvexity of the training procedure. When $d_{\mathrm{KL}}=1000$, the problem is essentially infinite-dimensional. The function space input FNMs (F2F and F2V) exhibit a nonparametric decay of test error as expected.
@@ 5.2. Aerodynamic force exerted on an airfoil
Here $\mathcal{A}$ denotes the closed curve defined by the union of the upper and lower surfaces of the airfoil. Different airfoil shapes are generated following the design element approach [19] (Figure 5). The initial NACA-0012 shape is embedded into a “cubic” design element featuring $8$ control nodes, and the initial shape is morphed to a different one following the displacement field of the control nodes of the design element. The displacements of control nodes are restricted to the vertical direction only. Consequently, the intrinsic dimension of the input is $7$, as displacing all nodes in the vertical direction by a constant value does not change the shape of the airfoil.
To generate the training data, we used the traditional second order finite volume method with the implicit backward Euler time integrator. The process begins by generating a new airfoil shape. Subsequently, a C-grid mesh [70] consisting of $221\times 51$ quadrilateral elements is created around the airfoil with adaptation near the airfoil. In total, we generated $2000$ training data and $400$ test data with the vertical displacements of each control node being sampled from a uniform distribution $\textsf{Uniform}([-0.05,0.05])$.
@@
Figure 7. Flow over an airfoil. Comparative analysis of data size versus relative test error for the FNM and NN approaches. The shaded regions denote two standard deviations away from the mean of the test error over $5$ realizations of the batch indices during SGD and model parameter initializations.
For each sample size $N$, five iid realizations of the models are trained on the relative loss (37) for $2000$ epochs in $L^{2}$ output space norm for functions and Euclidean norm for vectors. All FNM models use $4$ hidden layers, $12$ modes per dimension, and a channel width of $128$. We compare these models to a standard fully-connected NN with $4$ layers and a hidden width of $128$. In the case of FNM models, we observe that learning in the 1D setting consistently outperforms the 2D setting across all sizes of training data. Therefore, we only present results for the 1D setting. Moreover, this set of architectural hyperparameters with a large channel width of $128$ in general outperforms other hyperparameter settings. Figure 7 reveals several trends. As the data volume $N$ increases, all error curves decay at an algebraic rate that is slightly faster than $N^{-1/2}$. This may be due to the small sample sizes considered (under $2000$ data pairs) or, especially since the training data is noise-free, could be evidence of a data-driven “superconvergence” effect similar to that observed for QoI computations in adjoint methods for PDEs [21]. Overall, emulating PtO maps by training models with finite-dimensional vectors as both input and output (V2V and NN) is more challenging for this problem than adopting function space variants (F2F, F2V, V2F). The standard finite-dimensional NN performs similarly to V2V.
@@ 5.3. Effective tensor for a multiscale elliptic equation
and $\chi\colon\mathbb{T}^{2}\to\mathbb{R}^{2}$ solves the cell problem
@@
The bottleneck step in obtaining the effective tensor $\overline{A}$, which is our QoI, is solving the cell problem (53). Learning the solution map $A\mapsto\chi$ in (53) corresponds to the F2F setting and is explored in detail in [7]. Alternately, one could learn the effective tensor $\overline{A}$ directly using the F2V-FNM architecture to approximate $A\mapsto\overline{A}$. Furthermore, though $A$ is a function from $\mathbb{T}^{2}$ to $\mathbb{R}^{2\times 2}_{{\rm sym},\succ 0}$, in certain cases it may have an exact finite vector parametrization. One example of this case is finite piecewise-constant Voronoi tessellations; $A$ takes constant values on a fixed number of cells, and the cell centers uniquely determine the Voronoi geometry. Denoting these parameters as $z\in\mathbb{R}^{d_{u}}$ for appropriate $d_{u}\in\mathbb{N}$, one could also learn the V2F map $z\to\chi$ or the V2V map $z\to\overline{A}$. In this experiment, we compare the error in the QoI $\overline{A}$ using all four methods. A visualization of the possible maps is shown in Figure 8. Since our example is in two spatial dimensions, the $5$ Voronoi cell centers have two components each. The symmetry of $A$ yields three degrees of freedom (DoF) on each Voronoi cell. Altogether, this yields $25$ parameters that comprise the finite vector input.
@@ 6. Conclusion
This paper proposes the Fourier Neural Mappings (FNMs) framework as an operator learning method for approximating parameter-to-observable (PtO) maps with finite-dimensional input or output spaces, or both. Universal approximation theorems demonstrate that FNMs are well-suited for this task. Of central interest is the setting in which the PtO map factorizes into a vector-valued quantity of interest (QoI) map composed with a forward operator mapping between two function spaces. For this setting, the paper introduces the end-to-end (EE) and full-field (FF) learning approaches. The (EE) approach directly estimates the PtO map from its own input–output pairs, while the (FF) approach estimates the forward map first and then plugs this estimator into the QoI. The main theoretical results of the paper establish sample complexity bounds for Bayesian nonparametric regression of linear functionals with the (EE) and (FF) methods. The analysis reveals useful insights into how the smoothness of the QoI influences data efficiency. In particular, (FF) is superior to (EE) for smooth QoIs in this setting. The situation reverses for QoIs of low regularity. Finally, the paper implements the FNM architectures for three nonlinear problems arising from environmental science, aerodynamics, and materials modeling. The numerical results support the linear theory and extend beyond it by revealing the supremacy of function space representations of the input space over analogous finite-dimensional parametrizations.
@@
Several avenues for future work remain open. One way to understand the data efficiency of the (EE) and (FF) learning approaches beyond the specific Bayesian linear estimators that this paper analyzes would involve the development of fundamental lower bounds. Besides a few recent works [27, 33, 37, 43], there has been little attention on minimax lower bounds and (statistical) optimality for operator learning. The statistical theory in the present paper fixes an infinite-dimensional input space and studies the influence of the output space (being either one-dimensional or infinite-dimensional). The derivation of similar insights to those in Subsection 4.3.2 for PtO map learning with vector-to-function estimators could help explain the strong influence of the input space observed in the numerical experiments from Section 5. The scalar input and infinite-dimensional output case is partially addressed by [64]. Furthermore, it remains to be seen whether the theory developed in the present paper for the linear functional setting can extend to certain classes of nonlinear maps and QoIs, perhaps by linearizing the maps in an appropriate manner. For a specific nonlinear map, relevant work in [65] studies the approximation of nonsmooth QoIs. On the practical side, it is of interest to further explore architectural improvements for the various FNMs and in particular whether the latent function space introduced by the vector-to-vector FNM (M-V2V) can actually lead to improved performance over standard finite-dimensional neural networks.
@@
Acknowledgments
The authors thank Kaushik Bhattacharya for useful discussions about learning functionals, Andrew M. Stuart for helpful remarks about the universal approximation theory, and Zachary Morrow for providing the code for the advection–diffusion equation solver. The computations presented in this paper were partially conducted on the Resnick High Performance Computing Center, a facility supported by the Resnick Sustainability Institute at the California Institute of Technology.
Appendix A Additional variants of the main theorems
This appendix contains two additional theorems that are more general than their counterparts in the main text. While the main theorems as presented in Subsection 4.3 are sufficient to convey the primary message of this paper, we present the extra theorems here for completeness, as the results may be of independent interest.
@@ B.1. Supporting approximation results for neural operators
We need the following two lemmas that are simple generalizations of the universal approximation theorem for FNOs [30, Theorem 9, p. 9] to the setting where only one of the input or output domain is the torus. These results may be extracted from the proof of [30, Theorem 9, p. 9].
@@ C.1. Subgaussian and subexponential distributions
A real-valued random variable $X$ is said to be subgaussian [74, Section 2.5] if for some $\sigma>0$ it satisfies the moment generating function bound
@@
We write $X\in\operatorname{SG}(\sigma^{2})$ when (61) holds and define the subgaussian norm of $X$ by
@@
A real-valued random variable $Z$ is said to be subexponential [74, Section 2.7] if for some $v>0$ and $b>0$ it satisfies the moment generating function bound
@@
In contrast to the subgaussian case, the moment generating function of a subexponential random variable need only exist in a neighborhood of the origin instead of everywhere on the real line. We write $Z\in\operatorname{SE}(v^{2},\alpha)$ when (63) holds and define subexponential norm by
@@
The final term inside parentheses on the right hand side equals the subgaussian norm (upon replacing $2p$ with $p$). This is the asserted upper bound. The lower bound follows from the inequality $(\operatorname{\mathbb{E}}\lvert X\rvert^{2p})^{1/2}\geq\operatorname{\mathbb{E}}\lvert X\rvert^{p}$. ∎

### v2

@@ 5.1. Moments of an advection–diffusion model
Our first model problem concerns a canonical advection–diffusion PDE in two spatial dimensions. This equation often arises in the environmental sciences and is useful for modeling the spread of passive tracers (e.g., contaminants, pollutants, aerosols), especially when the driving velocity field is coupled to another PDE such as the Navier–Stokes equation. Our setup is as follows. Let $\mathcal{D}=(0,1)^{2}$ be the spatial domain and $\mathsf{n}$ denote the unit inward normal vector to $\mathcal{D}$. For a prescribed time-independent velocity field $v\colon\mathcal{D}\to\mathbb{R}^{2}$, the state $\phi\colon\mathcal{D}\times\mathbb{R}_{>0}\to\mathbb{R}$ solves
@@
for all $x=(x_{1},x_{2})\in\mathcal{D}$. Note that $u$ is constant in the vertical $x_{2}$ direction. The eigenvalues $\{\tau_{j}\}_{j\in\mathbb{N}}$ and eigenfunctions $\{e_{j}\}_{j\in\mathbb{N}}$ correspond to the Mercer decomposition of a kernel obtained by restricting a Matérn covariance function over $\mathbb{R}$ to $(0,1)\subset\mathbb{R}$. The covariance function has smoothness exponent $1.5$ and lengthscale $0.25$ [72]. We choose
$z_{j}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathsf{Uniform}([-1,1])\quad\text{for all}\quad j\in[d_{\mathrm{KL}}]\,.$
Thus, up to normalization constants, the velocity field (48) is the (truncated) KL expansion of a subgaussian stochastic process. We take the input to either be the full $x_{1}$-velocity field $u\colon\mathcal{D}\to\mathbb{R}$ or the i.i.d. realizations $z\coloneqq(z_{1},\ldots,z_{d_{\mathrm{KL}}})^{\top}$ of the random variables that affinely parametrize $u$.
Velocity Input State Output
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
@@
Define the nonlinear QoI map $q^{\dagger}\colon L^{4}(\mathcal{D};\mathbb{R})\to\mathbb{R}^{4}$ as follows. First, for any $h\in L^{2}(\mathcal{D};\mathbb{R})$, let
@@
Hence, $q^{\dagger}_{1}$ is the mean, $q^{\dagger}_{2}$ the standard deviation, $q^{\dagger}_{3}$ the skewness, and $q^{\dagger}_{4}$ the excess kurtosis. Our goal is to build FNM surrogates for the PtO map that sends the input representation (either the full velocity field or its finite number of i.i.d. coefficients) to the QoI values of the state $\phi$ at final time $t=3/4$ (see Figure 4). Therefore, we train FNMs to approximate each of the following ground truth maps:
@@
The training data is obtained by solving (47) with a second-order Lagrange finite element method on a mesh of size $32\times 32$ and Euler time step $0.01$. For each $d_{\mathrm{KL}}\in\{2,20,1000\}$, we generate $10^{4}$ i.i.d. data pairs for training, $1500$ pairs for computing the test error (which is (45) over the $1500$ test pairs instead of over the $N$ training pairs), and $500$ pairs for validation. All FNM models with 2D spatial input or output functions use $12$ modes per dimension and a channel width of $32$. For the V2V-FNM, we use a 1D latent function space with $12$ modes and channel width of $96$. We compare all FNM models to a standard fully-connected NN with $3$ layers and constant hidden width $2048$. These architecture settings were selected based on a hyperparameter search over the validation dataset for $d_{\mathrm{KL}}=1000$ that mimics the parameter complexity experiments in [8, 40]. The models are trained on the relative loss (45) for $500$ epochs in $L^{2}$ output space norm for functions and Euclidean norm for vectors. The optimizer settings include a minibatch size of $20$, weight decay of $10^{-4}$, and an initial learning rate of $10^{-3}$ which is halved every $100$ epochs. We train $5$ i.i.d. realizations of the models for various values of $N$ and $d_{\mathrm{KL}}$ and report the results in Figure 5.
@@
Figure 5 reveals several interesting trends. In general, training models to emulate the advection–diffusion PtO map with finite-dimensional vectors as input is more difficult than adopting function space input variants of the problem. The difficulty is further exacerbated as the dimension of the input vector (here, $d_{\mathrm{KL}}$) increases. We hypothesis that this gap in performance would reduce if the vector input models received the weighted KL coefficients $\{\sqrt{\tau_{j}}z_{j}\}$ as input instead of the i.i.d. sequence $\{z_{j}\}$. This way the model would have access to decay information and hence an ordering of the coefficients. The standard finite-dimensional NN performs poorly across all KL expansion dimensions. The training of the NN is also quite erratic, as evidenced by the large green shaded regions indicating large variance over multiple training runs. The output space seems to play less of a role than the input space. Indeed, the F2F and F2V FNMs with function space inputs generally achieve the lowest test error regardless of $N$ and $d_{\mathrm{KL}}$. The full-field F2F method slightly out performs the end-to-end F2V method by a small constant factor (except for when $d_{\mathrm{KL}}=2$). Since $q^{\dagger}$ is a smoothing QoI due to its integral definition, this observation aligns with the theoretical insights from Subsection 4.3.2. The fast convergence of some of the FNM models, especially for the low-dimensional cases $d_{\mathrm{KL}}=2$ and $d_{\mathrm{KL}}=20$, could potentially be explained by the lack of noise in the data, the smoothness of the QoIs, and the nonconvexity of the training procedure. When $d_{\mathrm{KL}}=1000$, the problem is essentially infinite-dimensional. The function space input FNMs (F2F and F2V) exhibit a nonparametric decay of test error as expected.
@@ 5.2. Aerodynamic force exerted on an airfoil
Here $\mathcal{A}$ denotes the closed curve defined by the union of the upper and lower surfaces of the airfoil. Different airfoil shapes are generated following the design element approach [21] (Figure 6). The initial NACA-0012 shape is embedded into a “cubic” design element featuring $8$ control nodes, and the initial shape is morphed into a different one following the displacement field of the control nodes of the design element. The displacements of control nodes are restricted to the vertical direction only. Consequently, the intrinsic dimension of the input is $7$, as displacing all nodes in the vertical direction by a constant value does not change the shape of the airfoil.
To generate the training data, we used the traditional second-order finite volume method with the implicit backward Euler time integrator. The process begins by generating a new airfoil shape. Subsequently, a $C$-grid mesh [81] consisting of $221\times 51$ quadrilateral elements is created around the airfoil with adaptation near the airfoil. In total, we generated $2000$ training data and $400$ test data with the vertical displacements of each control node being sampled from a uniform distribution $\textsf{Uniform}([-0.05,0.05])$.
@@
For each sample size $N$, five i.i.d. realizations of the models are trained on the relative loss (45) for $2000$ epochs in $L^{2}$ output space norm for functions and Euclidean norm for vectors. All FNM models use $4$ hidden layers, $12$ modes per dimension, and a channel width of $128$. We compare these models to a standard fully-connected NN with $4$ layers and a hidden width of $128$. In the case of FNM models, we observe that learning in the 1D setting consistently outperforms the 2D setting across all sizes of training data. Therefore, we only present results for the 1D setting. Moreover, this set of architectural hyperparameters with a large channel width of $128$ in general outperforms other hyperparameter settings.
Figure 8. Flow over an airfoil. Comparative analysis of relative test error versus data size for the FNM and NN approaches. The shaded regions denote two standard deviations away from the mean of the test error over $5$ realizations of the batch indices during SGD and model parameter initializations.
Figure 8 contains the results and reveals several trends. As the data volume $N$ increases, all error curves decay at an algebraic rate that is slightly faster than $N^{-1/2}$. This may be due to the small sample sizes considered (under $2000$ data pairs) or, especially since the training data is noise-free, could be evidence of a data-driven “superconvergence” effect similar to that observed for QoI computations in adjoint methods for PDEs [23]. Overall, emulating PtO maps by training models with finite-dimensional vectors as both input and output (V2V and NN) is more challenging for this problem than adopting function space variants (F2F, F2V, V2F). The standard finite-dimensional NN performs similarly to V2V.
@@ 5.3. Effective tensor for a multiscale elliptic equation
[…]
@@
[…]
@@ 6. Conclusion
This paper proposes the Fourier Neural Mappings (FNMs) framework as an operator learning method for approximating parameter-to-observable (PtO) maps with finite-dimensional vector inputs or outputs, or both. Universal approximation theorems demonstrate that FNMs are well-suited for this task. Of central interest is the setting in which the PtO map factorizes into a vector-valued quantity of interest (QoI) map composed with a forward operator mapping between two function spaces. For this setting, the paper introduces the end-to-end (EE) and full-field (FF) learning approaches. The (EE) approach directly estimates the PtO map from its own input–output pairs, while the (FF) approach estimates the forward map first and then plugs this estimator into the QoI. The main theoretical results of the paper establish sample complexity bounds for Bayesian nonparametric regression of linear functionals with the (EE) and (FF) methods. The analysis reveals useful insights into how the smoothness of the QoI influences data efficiency. In particular, (FF) is superior to (EE) for smooth QoIs in this setting. The situation reverses for QoIs of low regularity. A simulation study verifies these findings. Finally, the paper implements the FNM architectures for three nonlinear problems arising from environmental science, aerodynamics, and materials modeling. The numerical results support the linear theory and extend beyond it by revealing the supremacy of function space representations of the input space over analogous finite-dimensional vector parametrizations.
@@
Several avenues for future work remain open. One way to understand the data efficiency of the (EE) and (FF) learning approaches beyond the specific Bayesian linear estimators that this paper analyzes would involve the development of fundamental lower bounds. Besides a few recent works [1, 31, 36, 39, 43, 50], there has been little attention on minimax lower bounds and (statistical) optimality for operator learning. The statistical theory in the present paper fixes an infinite-dimensional input space and studies the influence of the output space (being either one-dimensional or infinite-dimensional). The derivation of similar insights to those in Subsection 4.3.2 for PtO map learning with vector-to-function estimators could help explain the strong influence of the input space observed in the numerical experiments from Section 5. The scalar input and infinite-dimensional output case is partially addressed by [75]. Furthermore, it remains to be seen whether the theory developed in the present paper for the linear functional setting can extend to certain classes of nonlinear maps and QoIs, perhaps by linearizing the maps in an appropriate manner. For a specific nonlinear map, relevant work in [76] studies the approximation of nonsmooth QoIs. On the practical side, it is of interest to further explore architectural improvements for the various FNMs and in particular whether the latent function space introduced by the vector-to-vector FNM (M-V2V) can actually lead to improved performance over standard finite-dimensional neural networks.
@@
Appendix A Additional variants of the sample complexity theorems
This appendix contains two additional theorems that are more general than their counterparts in Section 4 of the main text. While the main theorems as presented in Subsection 4.3 are sufficient to convey the primary message of that part of this paper, we present the extra theorems here for completeness, as the results may be of independent interest.
@@ B.1. Supporting approximation results for neural operators
We need the following two lemmas that are simple generalizations of the universal approximation theorem for FNOs [35, Theorem 9, p. 9] to the setting where only one of the input or output domain is the torus. These results may be extracted from the proof of [35, Theorem 9, p. 9].
@@ C.1. Subgaussian and subexponential distributions
A real-valued random variable $X$ is said to be subgaussian [86, Section 2.5] if for some $\sigma>0$ it satisfies the moment generating function bound
@@
We write $X\in\operatorname{\mathsf{SG}}(\sigma^{2})$ when (66) holds and define the subgaussian norm of $X$ by
@@
A real-valued random variable $Z$ is said to be subexponential [86, Section 2.7] if for some $v>0$ and $b>0$ it satisfies the moment generating function bound
@@
In contrast to the subgaussian case, the moment generating function of a subexponential random variable need only exist in a neighborhood of the origin instead of everywhere on the real line. We write $Z\in\operatorname{\mathsf{SE}}(v^{2},\alpha)$ when (68) holds and define subexponential norm by
@@
The final term inside parentheses on the right-hand side equals the subgaussian norm (upon replacing $2p$ with $p$). This is the asserted upper bound. The lower bound follows from the inequality $(\operatorname{\mathbb{E}}\lvert X\rvert^{2p})^{1/2}\geq\operatorname{\mathbb{E}}\lvert X\rvert^{p}$. ∎